# Strict comparison in reduced group $C^{*}$-algebras

### v2

@@ Abstract.
We prove that for every $n\geq 2$, the reduced group $C^{*}$-algebras of the countable free groups $C^{*}_{r}(\mathbb{F}_{n})$ have strict comparison. Our method works in a general setting: for $G$ in a large family of non-amenable groups, including hyperbolic groups, free products, mapping class groups, right-angled Artin groups etc., we have $C^{*}_{r}(G)$ have strict comparison. This work also has several applications in the theory of $C^{*}$-algebras including: resolving Leonel Robert’s selflessness problem for $C^{*}_{r}(G)$; uniqueness of embeddings of the Jiang-Su algebra $\mathcal{Z}$ up to approximate unitary equivalence into $C^{*}_{r}(G)$; full computations of the Cuntz semigroup of $C^{*}_{r}(G)$ and future directions in the $C^{*}$-classification program.
@@ 1.1. Strict comparison
The subspaces of a finite dimensional Hilbert space ${\mathcal{H}}$ can be compared and classified by their dimension; algebraically, the dimension of a subspace ${\mathcal{K}}\subset{\mathcal{H}}$ is the trace of the orthogonal projection onto ${\mathcal{K}}$. For matrices $x,y$ one has the following natural notion of comparison, that is $x\precsim y$ if $x=rys$ for some matrices $r,s$. It turns out here that the rank is what precisely implements this comparison, that is $\mathrm{rank}(x)\leq\mathrm{rank}(y)$ if and only if $x\precsim y$. Murray and von Neumann studied this comparison quite effectively in a broad setting of bounded operators on Hilbert spaces, in what are today called von Neumann algebras. In a broad class of von Neumann algebras known as II1 factors, they were able to establish a comparison theorem, showing that the lattice of projections under the above subequivalence is determined exactly by the unique finite trace [47]. Notably, in this context the trace takes on continuous values.
@@
Let $A$ be a $C^{*}$-algebra and let $\mathbb{K}$ denote the compact operators on $\ell^{2}$. If $a,b\in(A\otimes{\mathbb{K}})_{+}$ are positive elements, we say that $a\precsim b$ (Cuntz subequivalent) if there are $v_{n}\in A\otimes{\mathbb{K}}$ such that $v_{n}^{*}bv_{n}\to a$. If $a\precsim b$ and $b\precsim a$, we say that $a\sim b$. For $C^{*}$-algebras sometimes their behavior only becomes apparent at higher matrix levels, for instance the Bott projection which can only be seen in the $A\otimes\mathbb{M}_{2}(\mathbb{C}))$. This is why stabilization with the compacts is necessary. Now given $\tau$ a trace on $A$, there is a canonical extension of $\tau$ on $A$ to $(A\otimes{\mathbb{K}})_{+}$ (denoted by $\tau$ again). The dimension function here with respect to $\tau$ is defined on $(A\otimes{\mathbb{K}})_{+}$ by $d_{\tau}(a):=\lim_{n}\tau(a^{\frac{1}{n}})$ (note that for positive operators in matrices, $d_{\tau}$ precisely gives the rank of the operator). If $a,b\in(A\otimes{\mathbb{K}})_{+}$ satisfy $a\precsim b$, then $d_{\tau}(a)\leq d_{\tau}(b)$. A unital, simple, exact $C^{*}$-algebra $A$ with unique trace $\tau$ has strict comparison if whenever $a,b\in(A\otimes{\mathbb{K}})_{+}$ are such that $d_{\tau}(a)<d_{\tau}(b)$, then $a\precsim b$. Note that all $C^{*}$-algebras considered in the main results of our paper will be unital, simple, exact and with unique trace, so that the above definition will suffice. In full generality, the definition will involve the space of quasitraces and is more subtle.
Despite being such a fundamental and natural property of a $C^{*}$-algebra, proving strict comparison is difficult and has far-reaching applications (please see subsection 1.3 where several applications arising from our present work are described). Strict comparison occupies an important place as a hypothesis in Elliott’s important classification program, which seeks to classify simple, nuclear $C^{*}$-algebras by $K$-theoretical and tracial data. This program has seen immense success in recent times and has resulted in a substantial body of work due to several mathematicians (see for instance [30, 31, 13, 12, 67, 65]). Note also the influential Toms-Winter conjecture predicts that three regularity properties of very different natures, namely topological (finite nuclear dimension), functional analytic ($\mathcal{Z}$-stability), and algebraic (strict comparison), are equivalent for a separable, unital, simple, non-elementary, nuclear $C^{*}$-algebra $A$.
@@ 1.2. Context and main results
A prototypical example of a reduced group $C^{*}$-algebra is the one arising from the countable free group $\mathbb{F}_{2}$. The study of this $C^{*}$-algebra has been extremely important for opening up several deep insights in the subject. Some examples of this include the following fundamental results: unique trace and simplicity [52] (see also the major works [39, 35] which were inspired by Powers’ insights), lacking non-trivial projections [50], $K$-theory computations [50], and showing stable rank 1 [22] in the sense of Rieffel [54]. Despite these important developments, the fundamental question of proving strict comparison for $C^{*}_{r}(\mathbb{F}_{2})$ has evaded experts. Within the community, it has been a widely-known open problem with initial motivation stemming from the work of Dykema-Rørdam in 1998 ([24, 25], see also [55, Proposition 6.3.2]), which addresses comparison in $C^{*}_{r}(\mathbb{F}_{\infty})$. Note that the ideas behind the proof for strict comparison for $C^{*}_{r}(\mathbb{F}_{\infty})$ cannot be readily used to prove this result for $C^{*}_{r}(\mathbb{F}_{2})$. For instance, the proof there uses the ability to find freely independent elements from any finite subsets, up to norm perturbations, which holds inside $C^{*}_{r}(\mathbb{F}_{\infty})$. However such behavior is not clear at all to witness inside $C^{*}_{r}(\mathbb{F}_{2})$. Strict comparison for $C^{*}_{r}({\mathbb{F}}_{2})$ has been explicitly stated as an open question in the work of Robert in 2012 [55] and in the recent substantial problem list in $C^{*}$-algebras [60], see Problems 86 and 88.
@@
Another key motivation for showing strict comparison is that it is crucial in fully computing the Cuntz semigroup, introduced by Cuntz in his fundamental investigation on the existence of traces on $C^{*}$-algebras [18]. Over several decades, Cuntz semigroup theory has evolved into a very mature field of research [27]. It has been a long-sought problem to obtain full computations of the Cuntz semigroup of reduced group $C^{*}$-algebras in a general sense and has been described in [27] as ‘‘may be out of reach’’. The problem of computing the Cuntz semigroup for $C^{*}_{r}(\mathbb{F}_{2})$ is even explicitly stated as Problem 16.4 therein.
We settle all of the problems described above in this paper. Our first main result is the following:
@@
For any $n\geq 2$, $C^{*}_{r}(\mathbb{F}_{n})$ has strict comparison. In fact, $C^{*}_{r}(\mathbb{F}_{n})$ is selfless in the sense of Robert [56].
Thanks to major developments [57, 45], we now know that for unital, separable, simple, unique trace, nuclear, infinite-dimensional $C^{*}$-algebras, strict comparison and $\mathcal{Z}$-stability are equivalent. Importantly, at the moment, outside the $\mathcal{Z}$-stable case there are only a few examples of strict comparison arising from groups, such as infinite reduced free products like $C^{*}_{r}(\mathbb{F}_{\infty})$ ([55, Proposition 6.3.2]). With the next main result, we greatly expand the list of known examples. To motivate this, we need to discuss the following notion. We say a group $G$ has the rapid decay property if, for every finite generating set $X$, there exists a polynomial $P$ such that for every element $a\in\mathbb{C}G$ supported on $B_{X}(r)$, $\left\|\lambda_{G}(a)\right\|\leq P(r)\|a\|_{2}.$ The rapid decay property (see the survey [58]) was first proved by Haagerup for the free groups [33]. It was then generalized to Gromov hyperbolic groups [38, 36], relatively hyperbolic groups with peripheral subgroups having rapid decay [58], cocompact lattices in $SL_{3}(\mathbb{R})$ or $SL_{3}(\mathbb{C})$ [42, 14], mapping class groups of surfaces [7], large type Artin groups [15], heirarchically hyperbolic groups [5], and more. Rapid decay is extremely influential and has been used to prove several important problems, including the Novikov conjecture [16], the Baum-Connes conjecture [43], and others including [32]. Before we state our result, we will need to recall one more important and influential notion from geometric group theory [49]. A group is said to be acylindrically hyperbolic if it admits a non-elementary acylindrical action on a Gromov hyperbolic metric space. We direct the reader to [48] for context, examples, and references. We are now ready to state our next result:
@@
Let $G$ be an arbitrary countable group that is acylindrically hyperbolic with trivial finite radical and, in addition, enjoys the rapid decay property. Then $C^{*}_{r}(G)$ has strict comparison. In fact, we have $C^{*}_{r}(G)$ is selfless in the sense of Robert [56].
Note that the trivial finite radical assumption in the above theorem is necessary due to the fact that these $C^{*}$-algebras will have unique trace as a consequence of being selfless (see also [11]). The following are some examples of groups (with an added trivial finite radical assumption) that fit into the above result (see [48, 20]): non-elementary hyperbolic groups; non-elementary relatively hyperbolic groups relative to peripheral subgroups with the rapid decay property, which includes free products of groups with the rapid decay property; mapping class groups of compact orientable surfaces [8]; graph products of rapid decay groups over graphs without a splitting [46]; heirarchically hyperbolic groups [6]. In light of [42], it is very timely and interesting to ask if our main results hold in the setting of higher rank lattices, especially cocompact lattices in $SL_{3}(\mathbb{R})$. Our proof strategy has no bearing in this setting, and therefore we leave it as an important open question arising from our work. Notably, even stable rank $1$ seems to be an open problem for reduced group $C^{*}$-algebras of higher rank lattices.
Importantly, also see Remark 3.9, which describes why the rapid decay assumption cannot be removed with our current approach. However, we point out that our result gives new examples of groups which do not have rapid decay property but whose reduced $C^{*}$-algebras are selfless. Indeed by Theorem A and Theorem 4.2 of [56], we have $C^{*}_{r}(\mathbb{F}_{2}*SL_{3}(\mathbb{Z}))$ is selfless, however $\mathbb{F}_{2}*SL_{3}(\mathbb{Z})$ does not have rapid decay.
@@ 1.3. Applications to classification and Cuntz semigroup theory
It is a very standard fact that every II1 factor contains a copy of the hyperfinite II1 factor $R$ uniquely up to pointwise approximate unitary equivalence. The most naïve analogue of this in $C^{*}$-algebras would be that every simple, unital, infinite-dimensional $C^{*}$-algebra contains the Jiang-Su algebra $\mathcal{Z}$ uniquely up to pointwise approximate unitary equivalence. Both existence and uniqueness problems for embeddings of $\mathcal{Z}$ tend to be difficult, see [60, Section 18]. It is known that there are simple, unital, infinite-dimensional $C^{*}$-algebras that do not contain $\mathcal{Z}$ [19]. It remains open if there is any unital $C^{*}$-algebra $A$ with two non-equivalent unital embedding of $\mathcal{Z}$. Our work yields the following concrete application. The result below follows from Proposition 6.3.1 in [55] and our Theorem B. Note that for our examples, the uniqueness of the trace follows from the fact that the groups in consideration have trivial amenable radical and the unique quasitrace comes from exactness and Haagerup’s famous quasitrace theorem [34]. Furthermore, stable rank one holds for all acylindrically hyperbolic groups [28]. Note also that the following result addresses problems 62 and 64 in [60]:
@@
For any group $G$ in the class of groups covered in Theorem B, there is a unital embedding of the Jiang-Su algebra $\mathcal{Z}\to C^{*}_{r}(G)$, and this embedding is unique up to pointwise approximate unitary equivalence.
@@
As explained in the beginning of the introduction, strict comparison is a crucial hypothesis that is used in classifying maps up to $K$-theoretic and tracial data in the nuclear setting. In the natural progression of the current landscape of classification, one expects new developments to emerge in the more general setting of non-nuclear $C^{*}$-algebras, wherein strict comparison will still naturally play a crucial role as a hypothesis for the co-domain for classifying maps between $C^{*}$-algebras up to the total invariant. Our results here will provide a wealth of interesting new and natural co-domains for this type of classification phenomena.
@@
In [17], Coward-Elliott-Ivanescu gave a new picture of the Cuntz semigroup in terms of Hilbert modules; they showed that for $C^{*}$-algebras of stable rank one, the Cuntz semigroup encodes precisely the isomorphism classes of countably generated Hilbert modules, similarly to how the Murray-von Neumann semigroup encodes isomorphism classes of finitely generated projective modules. The theory of Cuntz semigroups is particularly well-developed for $C^{*}$-algebras of stable rank one [1]. An important notion in this context is pureness, which was introduced by Winter in his seminal study [66] of regularity properties for simple, nuclear $C^{*}$-algebras. Pureness can be thought of as $\mathcal{Z}$-stability at the level of the Cuntz semigroup, since in [2], it was shown that a $C^{*}$-algebra $A$ is pure if and only if its Cuntz semigroup $\mathrm{Cu}(A)$ tensorially absorbs $\mathrm{Cu}(\mathcal{Z})$. More recently, it was shown that pure $C^{*}$-algebras form a robust class [3], and as observed in [3, Example 5.14], if $G$ is an exact, infinite, $C^{*}$-simple group, then $C^{*}_{r}(G)$ has strict comparison if and only if $C^{*}_{r}(G)$ is pure. If $A$ is a unital, separable, simple, stably finite, pure $C^{*}$-algebra, then the Cuntz semigroup can be computed as $\mathrm{Cu}(A)=V(A)\sqcup\mathrm{LAff}(QT(A))_{++}$, where $QT(A)$ denotes the quasitracial states on $A$. If $A$ is exact, then $QT(A)=T(A)$ by a famous result of Haagerup [34]. Thus, if G is an exact, infinite, C*-simple group such that $C^{*}_{r}(G)$ has strict comparison, then its Cuntz semigroup can be computed as $\mathrm{Cu}(C^{*}_{r}(G))=V(C^{*}_{r}(G))\sqcup[0,\infty]$ since such $C^{*}$-algebras have a unique trace [11]. Combined with our main results, this answers Problem 16.4 from [27] which asks to fully compute the Cuntz semigroup for $C^{*}_{r}(\mathbb{F}_{n})$, which is a Cuntz semigroup version of Blackadar’s fundamental comparison property [10] for projections in $C^{*}_{r}(\mathbb{F}_{n})$.
@@
The Cuntz semigroup of $C^{*}_{r}(\mathbb{F}_{2})$ is $\mathbb{N}\sqcup[0,\infty]$ with the natural semigroup structure as in $\mathcal{Z}$.
@@
Note that, our Theorem B can be used to fully compute the Cuntz semigroup for the vast family of acylindrically hyperbolic groups with trivial finite radical and with rapid decap property as described above. This partially settles the general problem in the paragraph before Problem 16.4 in [27], obtaining a full computation for many classes of $C^{*}$-simple groups. For examples of groups with stable rank one, see [22, 28, 53]. Note also that Theorem B additionally recovers with a new proof stable rank one for the families of groups in the statement (via [56]).
1.4. Functional analytic Tarski problem
We mention here that Theorem A in this paper will play a key role in the upcoming joint work [59] of the third author and C. Schafhauser, obtaining a negative resolution to Tarski problem on the first order theory of reduced free group $C^{*}$-algebras, showing that they are pairwise non elementarily equivalent. This is in sharp contrast to the positive solution of Tarski’s problem on the first order theory of free groups, obtained in the remarkable independent works of Sela and Kharlampovich-Myasnikov [61, 40].
1.5. Comments

### v4

@@ Abstract.
We prove that for every $n\in\mathbb{N}$ such that $n\geq 2$, the reduced group $C^{*}$-algebras of the countable free groups $C^{*}_{r}(\mathbb{F}_{n})$ have strict comparison. Our method works in a general setting: for every finitely generated acylindrically hyperbolic group $G$ with trivial finite radical and the rapid decay property, we have $C^{*}_{r}(G)$ have strict comparison. This work also has several applications in the theory of $C^{*}$-algebras including: resolving Leonel Robert’s selflessness problem for $C^{*}_{r}(G)$; uniqueness of embeddings of the Jiang-Su algebra $\mathcal{Z}$ up to approximate unitary equivalence into $C^{*}_{r}(G)$; full computations of the Cuntz semigroup of $C^{*}_{r}(G)$ and future directions in the $C^{*}$-classification program.
@@ 1.1. Strict comparison
The subspaces of a finite dimensional Hilbert space $\mathcal{H}$} can be compared and classified by their dimension; algebraically, the dimension of a subspace $\mathcal{K}$$\subset{\mathcal{H}}$ is the trace of the orthogonal projection onto $\mathcal{K}$}. For matrices $x,y$ one has the following natural notion of comparison, that is $x\precsim y$ if $x=rys$ for some matrices $r,s$. It turns out here that the rank is what precisely implements this comparison, that is $\mathrm{rank}(x)\leq\mathrm{rank}(y)$ if and only if $x\precsim y$. Murray and von Neumann studied this comparison quite effectively in a broad setting of bounded operators on Hilbert spaces, in what are today called von Neumann algebras. In a broad class of von Neumann algebras known as II1 factors, they were able to establish a comparison theorem, showing that the lattice of projections under the above subequivalence is determined exactly by the unique finite trace [49]. Notably, in this context the trace takes on continuous values, that is, for all $t\in[0,1]$, there is a projection $p\in\mathcal{P}(M)$ with $\tau(p)=t$.
@@
Let $A$ be a $C^{*}$-algebra and let $\mathbb{K}$ denote the compact operators on $\ell^{2}$. If $a,b\in(A\otimes{\mathbb{K}})_{+}$ are positive elements, we say that $a\precsim b$ (Cuntz subequivalent) if there are $v_{n}\in A\otimes{\mathbb{K}}$ such that $v_{n}^{*}bv_{n}\to a$. Now given $\tau$ a trace on $A$, there is a canonical extension of $\tau$ to $(A\otimes{\mathbb{K}})_{+}$ (denoted by $\tau$ again). The dimension function here with respect to $\tau$ is defined on $(A\otimes{\mathbb{K}})_{+}$ by $d_{\tau}(a):=\lim_{n}\tau(a^{\frac{1}{n}})$ (note that for positive operators in matrices, $d_{\tau}$ precisely gives the rank of the operator). If $a,b\in(A\otimes{\mathbb{K}})_{+}$ satisfy $a\precsim b$, then $d_{\tau}(a)\leq d_{\tau}(b)$. A unital, simple, exact $C^{*}$-algebra $A$ with unique trace $\tau$ has strict comparison if whenever $a,b\in(A\otimes{\mathbb{K}})_{+}$ are such that $d_{\tau}(a)<d_{\tau}(b)$, then $a\precsim b$. Note that many of the $C^{*}$-algebras considered in the results of our paper will be unital, simple, exact and with unique trace (this includes the reduced $C^{*}$-algebras of hyperbolic groups [1]), so that the above definition will suffice. In full generality, the definition will involve the space of quasitraces and is more subtle.
Despite being such a fundamental and natural property of a $C^{*}$-algebra, proving strict comparison is difficult and has far-reaching applications (please see subsection 1.3 where several applications arising from our present work are described). Strict comparison occupies an important place as a hypothesis in Elliott’s classification program, which seeks to classify simple, nuclear $C^{*}$-algebras by $K$-theoretical and tracial data. This program has seen immense success in recent times and has resulted in a substantial body of work due to several mathematicians (see for instance [32, 33, 15, 14, 71, 69]). Note also the influential Toms-Winter conjecture predicts that three regularity properties of very different natures, namely topological (finite nuclear dimension), functional analytic ($\mathcal{Z}$-stability), and algebraic (strict comparison), are equivalent for a separable, unital, simple, non-elementary, nuclear $C^{*}$-algebra $A$. These three seemingly independent regularity properties, which are conjectured to be equivalent, are exactly the conditions under which one considers a $C^{*}$-algebra to belong to the classification program.
@@ 1.2. Context and main results
A prototypical example of a reduced group $C^{*}$-algebra is the one arising from the countable free group $\mathbb{F}_{2}$. The study of this $C^{*}$-algebra has been extremely important for opening up several deep insights in the subject. Some examples of this include the following fundamental results: unique trace and simplicity [55] (see also the major works [42, 37] which were inspired by Powers’ insights), lacking non-trivial projections [53], $K$-theory computations [53], showing stable rank 1 [25] in the sense of Rieffel [57], and showing strong convergence matrix models [38] (see also [46] whose ideas inspired the present paper). Despite these important developments, the fundamental question of proving strict comparison for $C^{*}_{r}(\mathbb{F}_{2})$ has evaded experts. Within the community, it has been a widely-known open problem with initial motivation stemming from the work of Dykema-Rørdam in 1998 ([27, 28], see also [58, Proposition 6.3.2]), which addresses comparison in $C^{*}_{r}(\mathbb{F}_{\infty})$. Note that the ideas behind the proof for strict comparison for $C^{*}_{r}(\mathbb{F}_{\infty})$ cannot be readily used to prove this result for $C^{*}_{r}(\mathbb{F}_{2})$. For instance, the proof there uses the ability to find freely independent elements from any finite subsets, up to norm perturbations, which holds inside $C^{*}_{r}(\mathbb{F}_{\infty})$. However such behavior is not clear at all to witness inside $C^{*}_{r}(\mathbb{F}_{2})$. Strict comparison for $C^{*}_{r}({\mathbb{F}}_{2})$ has been explicitly stated as an open question in the work of Robert in 2012 [58] and in the recent substantial problem list in $C^{*}$-algebras [62], see Problems 86 and 88.
@@
Another key motivation for showing strict comparison is that it is crucial in fully computing the Cuntz semigroup, introduced by Cuntz in his fundamental investigation on the existence of traces on $C^{*}$-algebras [20]. Over several decades, Cuntz semigroup theory has evolved into a very mature field of research [29]. It has been a long-sought problem to obtain full computations of the Cuntz semigroup of reduced group $C^{*}$-algebras in a general sense and has been described in [29] as ‘‘may be out of reach’’. The problem of computing the Cuntz semigroup for $C^{*}_{r}(\mathbb{F}_{2})$ is even explicitly stated as Problem 16.4 therein. All of the problems described above are settled in this paper.
@@
For any $n\in\mathbb{N}$ such that $n\geq 2$, $C^{*}_{r}(\mathbb{F}_{n})$ has strict comparison. In fact, $C^{*}_{r}(\mathbb{F}_{n})$ is selfless in the sense of Robert [59].
Thanks to major developments [60, 47], we now know that for unital, separable, simple, unique trace, nuclear, infinite-dimensional $C^{*}$-algebras, strict comparison and $\mathcal{Z}$-stability are equivalent. Importantly, at the moment, outside the $\mathcal{Z}$-stable case there are only a few examples of strict comparison arising from groups, such as infinite reduced free products like $C^{*}_{r}(\mathbb{F}_{\infty})$ ([58, Proposition 6.3.2]). With our next main result, we greatly expand the list of known examples. To motivate this, we need to discuss the following notion. We say a group $G$ has the rapid decay property if, for every finite generating set $X$, there exists a polynomial $P$ such that for every element $a\in\mathbb{C}G$ supported on $B_{X}(r)$, $\left\|\lambda_{G}(a)\right\|\leq P(r)\|a\|_{2}.$ The rapid decay property (see the survey [61]) was first proved by Haagerup for the free groups [35]. It was then generalized to Gromov hyperbolic groups [41, 39], relatively hyperbolic groups with peripheral subgroups having rapid decay [61], cocompact lattices in $SL_{3}(\mathbb{R})$ or $SL_{3}(\mathbb{C})$ [44, 16], mapping class groups of surfaces [8], large type Artin groups [17], heirarchically hyperbolic groups [6], and more. Rapid decay is extremely influential and has been used to prove several important problems, including the Novikov conjecture [18], the Baum-Connes conjecture [45], and others including [34]. Before we state our result, we will need to recall one more important and influential notion from geometric group theory [51]. A group is said to be acylindrically hyperbolic if it admits a non-elementary acylindrical action on a Gromov hyperbolic metric space. We direct the reader to [50] for context, examples, and references. Recall also the finite radical for a group is the largest finite normal subgroup. Within the class of acylindrically hyperbolic groups $G$, it turns out that the finite radical being trivial is equivalent to $C^{*}_{r}(G)$ having unique trace, which is in turn equivalent to $C^{*}_{r}(G)$ being simple (see Theorem 1.4 of [12] and Theorem 2.35 of [22]). We are now ready to state our next result:
@@
Let $G$ be an arbitrary finitely generated countable group that is acylindrically hyperbolic with trivial finite radical and, in addition, enjoys the rapid decay property. Then $C^{*}_{r}(G)$ has strict comparison. In fact, we have $C^{*}_{r}(G)$ is selfless in the sense of Robert [59].
Note that the trivial finite radical assumption in the above theorem is necessary due to the fact that these $C^{*}$-algebras will have unique trace as a consequence of being selfless (see also [12]). The following are some examples of groups (with an added trivial finite radical assumption) that fit into the above result (see [50, 22]): non-elementary hyperbolic groups; non-elementary relatively hyperbolic groups relative to peripheral subgroups with the rapid decay property, which includes free products of groups with the rapid decay property; mapping class groups of compact orientable surfaces [9]; graph products of rapid decay groups over graphs without a splitting [48]; heirarchically hyperbolic groups [7]. In light of [44], it is very timely and interesting to ask if our main results hold in the setting of higher rank lattices, especially cocompact lattices in $SL_{3}(\mathbb{R})$  [^1] In a recent development [65], I. Vigdorovich has settled this problem.. Our proof strategy has no bearing in this setting, and therefore we leave it as an important open question arising from our work. Notably, even stable rank $1$ seems to be an open problem for reduced group $C^{*}$-algebras of higher rank lattices. We also speculate that our proof technique might open up a way to prove strict comparison for even wider families of groups, such as all $C^{*}$-simple groups with rapid decay, see Remark 3.6.  [^2] In a very recent development [52], N. Ozawa has developed a new approach that proves selflessness for the reduced $C^{*}$-algebras of groups admitting extremely proximal boundaries.
Importantly, also see Remark 3.9, which describes why the rapid decay assumption cannot be removed with our current approach. However, we point out that our result gives new examples of groups which do not have rapid decay property but whose reduced $C^{*}$-algebras are selfless. Indeed by Theorem A and Theorem 4.2 of [59], we have $C^{*}_{r}(\mathbb{F}_{2}*SL_{3}(\mathbb{Z}))$ is selfless, however $\mathbb{F}_{2}*SL_{3}(\mathbb{Z})$ does not have rapid decay (see Remark 8.6 in [64]).
@@ 1.3. Applications to classification and Cuntz semigroup theory
It is a very standard fact that every II1 factor contains a copy of the hyperfinite II1 factor $R$ uniquely up to pointwise approximate unitary equivalence. The most naïve analogue of this in $C^{*}$-algebras would be that every simple, unital, infinite-dimensional $C^{*}$-algebra contains the Jiang-Su algebra $\mathcal{Z}$ uniquely up to pointwise approximate unitary equivalence. Both existence and uniqueness problems for embeddings of $\mathcal{Z}$ tend to be difficult, see [62, Section 18]. It is known that there are simple, unital, infinite-dimensional $C^{*}$-algebras that do not contain $\mathcal{Z}$ [21]. It remains open if there is any unital $C^{*}$-algebra $A$ with two non-equivalent unital embedding of $\mathcal{Z}$. Our work yields the following concrete application. The result below follows from Proposition 6.3.1 in [58] (which can be applied since the reduced $C^{*}$-algebras of the groups in consideration have unique quasitrace from Haagerup’s famous quasitrace theorem [36], and have stable rank one from [30] see also [56]) and our Theorem B. Note also that the following result addresses problems 62 and 64 in [62]:
@@
For any exact group $G$ in the class of groups covered in Theorem B, there is a unital embedding of the Jiang-Su algebra $\mathcal{Z}\to C^{*}_{r}(G)$, and this embedding is unique up to pointwise approximate unitary equivalence.
@@
As explained in the beginning of the introduction, strict comparison is a crucial hypothesis that is used in classifying maps up to $K$-theoretic and tracial data in the nuclear setting. In the natural progression of the current landscape of classification, we expect new developments to emerge in the more general setting of non-nuclear $C^{*}$-algebras, wherein strict comparison will still naturally play a crucial role as a hypothesis for the co-domain for classifying maps between $C^{*}$-algebras up to the total invariant. Our results here will provide a wealth of interesting new and natural co-domains for this type of classification phenomena.
@@
In [19], Coward-Elliott-Ivanescu gave a new picture of the Cuntz semigroup in terms of Hilbert modules; they showed that for $C^{*}$-algebras of stable rank one, the Cuntz semigroup encodes precisely the isomorphism classes of countably generated Hilbert modules, similarly to how the Murray-von Neumann semigroup encodes isomorphism classes of finitely generated projective modules. The theory of Cuntz semigroups is particularly well-developed for $C^{*}$-algebras of stable rank one [2]. An important notion in this context is pureness, which was introduced by Winter in his seminal study [70] of regularity properties for simple, nuclear $C^{*}$-algebras. Pureness can be thought of as $\mathcal{Z}$-stability at the level of the Cuntz semigroup, since in [3], it was shown that a $C^{*}$-algebra $A$ is pure if and only if its Cuntz semigroup $\mathrm{Cu}(A)$ tensorially absorbs $\mathrm{Cu}(\mathcal{Z})$. More recently, it was shown that pure $C^{*}$-algebras form a robust class [4], and as observed in [4, Example 5.14], if $G$ is an exact, infinite, $C^{*}$-simple group, then $C^{*}_{r}(G)$ has strict comparison if and only if $C^{*}_{r}(G)$ is pure. If $A$ is a unital, separable, simple, stably finite, pure $C^{*}$-algebra, then the Cuntz semigroup can be computed as $\mathrm{Cu}(A)=V(A)\sqcup\mathrm{LAff}(QT(A))_{++}$ (see Theorem 2.6 in [13]). Here where $QT(A)$ denotes the quasitracial states on $A$, $V(A)$ denotes the Murray-von Neumann semigroup and $\mathrm{LAff}(K)$ means the lower-semicontinuous, affine functions $K$ to $(-\infty,\infty]$ if $K$ is a compact convex set, and $\mathrm{LAff}(K)_{++}$ means the subset of strictly positive elements, that is, the lower-semicontinuous, affine functions $K$ to $(0,\infty]$. If $K$ contains only one point, then $\mathrm{LAff}(K)_{++}=(0,\infty]$. If $A$ is exact, then $QT(A)=T(A)$ by a famous result of Haagerup [36]. Thus, if G is an exact, infinite, C*-simple group such that $C^{*}_{r}(G)$ has strict comparison, then its Cuntz semigroup can be computed as $\mathrm{Cu}(C^{*}_{r}(G))=V(C^{*}_{r}(G))\sqcup[0,\infty]$ since such $C^{*}$-algebras have a unique trace [12]. Combined with our main results, this answers Problem 16.4 from [29] which asks to fully compute the Cuntz semigroup for $C^{*}_{r}(\mathbb{F}_{n})$, which is a Cuntz semigroup version of Blackadar’s fundamental comparison property [11] for projections in $C^{*}_{r}(\mathbb{F}_{n})$.
@@
The Cuntz semigroup of $C^{*}_{r}(\mathbb{F}_{2})$ is $\mathbb{N}\sqcup[0,\infty]$.
@@
Note that, our Theorem B can be used to fully compute the Cuntz semigroup for the vast family of exact acylindrically hyperbolic groups with trivial finite radical and with rapid decay property as described above. This partially settles the general problem in the paragraph before Problem 16.4 in [29], obtaining a full computation for many classes of $C^{*}$-simple groups. For examples of groups with stable rank one, see [25, 30, 56]. Note our work also recovers with a new proof, stable rank one for the families of groups in the statement of Theorem B (via Theorem 3.1(2) of [59]).
1.4. Comments to the reader